OpenStudy (anonymous):

Sequence formula for {3,7,3,7..}

OpenStudy (anonymous):

Find a general formula for this set.

OpenStudy (freckles):

you have the odd terms are constantly 3 and the even terms are constantly 7 set up a piecewise function saying exactly that

OpenStudy (anonymous):

I don't think they will allow piecewise functions on my homework.

OpenStudy (anonymous):

Or is there any other way to figure this out??

OpenStudy (freckles):

well I suppose we can write it another way so we have (-1)^n=1 if n is odd and -1 if n is even I think we can use this to get our 3 and out 7 for example pretend we have 5,9,5,9,... so we can do this \[a_n=\frac{9+5}{2}+\frac{9-5}{2}(-1)^n \\ a_n=7+2(-1)^n \]

OpenStudy (anonymous):

Wow, that is VERY smart

OpenStudy (freckles):

I think that will always work for sequences like this a,b,a,b,... assume b>a \[A_n=\frac{b+a}{2}+\frac{b-a}{2}(-1)^n\]

OpenStudy (freckles):

I called the sequence big A so not confuse with the little number a in that sequence (remember math is case sensitive) :)

OpenStudy (anonymous):

but

OpenStudy (anonymous):

OH never mind (: Thank you

OpenStudy (anonymous):

Is this a formula I should memorize?